OpenStudy (anonymous):

scott earns $5 for each lawn he mows.on saturday he mowed 6 lawns,and on sunday he mowed 8 lawns.which of the following expressions represents how much scott earned for the weekend ( in dollars)

OpenStudy (anonymous):

6 * 5 = 30. (For Saturday) 6 * 8 = 48 (For Sunday)